OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). (2 points) y = 2x – 7 y = −x − 2 y = 2x – 2 y = −x − 7 2. Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). (2 points) y = 2x + 6 y = −x + 3 y = −x + 6 y = 2x + 3 3. Write the equation of the line that is parallel to the line 4x – 3y = −12 and passes through the point (−3, 4). (2 points) y = x + 8 y = x + 3 y = −x + 8 y = −x + 3

OpenStudy (anonymous):

4. Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, –1). (2 points) y = x − 3 y = x + 17 y = –3x − 3 y = –3x + 17 5. Explain, in complete sentences, the relationship between the equations of parallel lines. (2 points)

OpenStudy (anonymous):

Those too please!!

OpenStudy (anonymous):

1. A 2. C (If the slope is -1/2x instead of -x) 3. D 4. A (If the slope is 1/3x instead of x) 5. Perpendicular lines have slopes that are opposite reciprocals of each other (different signs, flipped fractions, i.e. 2/3, -3/2)

OpenStudy (anonymous):

Thankyou so much! @KirstiSmith

OpenStudy (anonymous):

Welcome!